# Surf-D: High-Quality Surface Generation for Arbitrary Topologies using Diffusion Models

### 28 Nov 2023 v1

@@ Abstract
In this paper, we present Surf-D, a novel method for generating high-quality 3D shapes as Surfaces with arbitrary topologies using Diffusion models. Specifically, we adopt Unsigned Distance Field (UDF) as the surface representation, as it excels in handling arbitrary topologies, enabling the generation of complex shapes. While the prior methods explored shape generation with different representations, they suffer from limited topologies and geometry details.
Moreover, it’s non-trivial to directly extend prior diffusion models to UDF because they lack spatial continuity due to the discrete volume structure. However, UDF requires accurate gradients for mesh extraction and learning. To tackle the issues, we first leverage a point-based auto-encoder to learn a compact latent space, which supports gradient querying for any input point through differentiation to effectively capture intricate geometry at a high resolution. Since the learning difficulty for various shapes can differ, a curriculum learning strategy is employed to efficiently embed various surfaces, enhancing the whole embedding process. With pretrained shape latent space, we employ a latent diffusion model to acquire the distribution of various shapes. Our approach demonstrates superior performance in shape generation across multiple modalities and conducts extensive experiments in unconditional generation, category conditional generation, 3D reconstruction from images, and text-to-shape tasks. Visit our project page at https://yzmblog.github.io/projects/SurfD/.
@@
11footnotetext: Equal Contributions.
1 Introduction
Crafting diverse 3D shapes for engaging virtual worlds has been challenging, requiring creativity and extensive design skills. Recent advancements for 3D shape generation [4, 1, 41, 5, 22, 34, 40] aim to make this process more accessible.
@@ 1 Introduction
The diverse interplay of topology and intricate geometry gives rise to countless 3D shapes in our world. However, existing generation works still suffer from limited topologies and the lack of geometric details.
@@
Previous methods present shapes using point clouds [1, 41], voxel representations [57, 71], and signed distance fields (SDFs) [11, 44]. However, point clouds and voxels encounter difficulties in faithfully extracting surfaces. Implicit representations, particularly SDFs, excel in describing closed surfaces but face challenges with complex topologies. Despite attempts to enhance them, e.g. 3PSDF [10] and TSDF [60], the representations remain reliant on the underlying definition of SDF. G-Shell [38] is proposed to represent open surfaces but still falters with non-manifold structures and computational constraints.
In this paper, we propose Surf-D, the first generative latent Diffusion model employing Unsigned Distance Fields (UDF) which is a general shape representation for any 3D shape, for versatile shape generation with high-quality surfaces and arbitrary topologies.
Surf-D first utilizes an AutoEncoder to compress 3D surfaces into a compact low-dimensional representation during shape embedding and learns a continuous distance field. However, producing high-resolution 3D surfaces with detailed geometry by UDF with the latent diffusion model is a non-trivial effort. Different from other shape representations, the continuous gradient field is essential to UDF’s learning [74] and meshing [23]. In contrast, previous volume-based generative frameworks [12, 20, 56] generate the discrete grid with SDF values, which is computationally costly owing to the cubic space complexity of their voxel grid and failed to access the accurate gradient field when applying to UDFs.
@@
To tackle this problem, we develop a point-based AutoEncoder for learning intricate geometry described by UDFs as low-dimensional global latent codes rather than a volume-based manner [12, 20, 56]. Our point-based AutoEncoder allows gradient querying for each input point through differentiation, contributing reliable pseudo-signs in bisection search [23] that are commonly used for surface extraction from the implicit fields and gradient field constrains [74] when learning UDFs. As a result, our model is not only much more lightweight and learning-efficient but also feasible to convert high-quality generated implicit shapes to explicit meshes. A comparison with volume-based shape embedding [12] with the same resolution setting is in Sec. 4.4. Additionally, to learn the compact shape latent space efficiently and effectively on diverse surfaces, we employ a curriculum learning strategy [2] to train our model in an easy-to-hard sampling order, i.e. progressively extending the current training dataset during training. This strategy promotes our model to develop a diverse latent space, avoiding entrapment into generating mean shape, i.e. model collapse.
@@
    We conduct latent diffusion and devise a point-based AutoEncoder that enables effective gradient querying for input points, achieving high-resolution and detail-preserving surface generation. To enhance surface learning, we implement a curriculum learning scheme.
@@ 2 Related Work
Tremendous efforts have been made for 3D shape generation [4, 1, 41, 5, 22, 37, 7], given its broad applications, employing various shape representations such as voxel grids [69, 67], point clouds [41, 63, 79], and implicit neural fields [47, 18, 12, 26]. Additionally, various methods have been developed to guide the synthesis with text prompt [54, 20, 39, 9, 36], single-view image [34, 40, 33, 32], sketch [70, 76] and sparse control anchors [42, 14]. In the early stages, voxel-based generation methods [69, 67, 50] can synthesize various coarse shapes. However, the voxel representation can not scale to high-resolution generation [78, 76], due to its cubic space complexity. Utilizing point clouds [41, 63, 79], which are compact and readily available, for 3D object representation can achieve arbitrary shape generation with their well-suitability to learn geometry and texture information [45, 70]. However, it requires complex post-processing to faithfully convert the synthesized point clouds into meshes for real-world applications [48]. Recent works employ implicit neural fields for modeling occupancy [19, 75], signed distance functions (SDF) [12, 76], and truncated signed distance (T-SDF) [20, 26], yielding impressive generation results with continuous surface. Despite these advancements, they struggle to accurately represent complex shapes with diverse topologies, e.g. non-watertight surfaces [74, 13]. Although the very recent work G-Shell [38] presents a method for open surface modeling, it encounters challenges with non-manifold structures and computational constraints. In contrast, our method achieves a generalized surface generation with arbitrary topologies using UDFs with latent diffusion for high-resolution 3D shape synthesis.
@@ 3.1.2 Loss Functions
3.1.3 Curriculum Learning Strategy
@@
Curriculum training involves measuring performance and establishing an easy-to-hard scheduler. In this paper, we use a simple yet effective performance metric, which is the training loss $\mathcal{L}_{latent}$. We select samples with lower losses as incoming training data in each iteration. This continues until all data is included. Specifically, we first randomly sample $k$ shapes from the dataset, then gradually add new difficult samples according to their loss values, resulting in an easy-to-head gradual learning process. The pseudo-code can be expressed as Alg. 1. The influence of the proposed learning strategy is included in the Appendix.
Algorithm 1 Curriculum Learning Strategy
0:  $M$ – our model;
0:  $E$ – the whole dataset;
0:  $E^{*}$ – current training dataset;
0:  $n$ – number of epochs;
0:  $s$ – number of switch epochs;
0:  $k$ – number of selected samples;
0:  $C$ – loss criterion
0:  $L$ – calculated losses
0:  $argkmin$ – return indexes of k minimal numbers
1:  $E^{*}\leftarrow random\_select(E,k)$
2:  for $t\in{1,2,...,n}$ do
3:     if ($t\%s$ == 00) and (not $E^{*}$ == $E$then
4:        $L\leftarrow C(M,E-E^{*})$
5:        $E^{*}\leftarrow select(E-E^{*},argkmin(L))$
6:     end if
7:     $M\leftarrow train(M,E^{*},C)$
8:  end for
@@ 3.2 Latent Diffusion Model for UDF
$\begin{split}q\left(\bm{z}_{0:T}\right)&=q\left(\bm{z}_{0}\right)\Pi_{t=1}^{T}q\left(\bm{z}_{t}|\bm{z}_{t-1}\right),\\ p_{\theta}\left(\bm{z}_{0:T}\right)&=p\left(\bm{z}_{T}\right)\Pi_{t=1}^{T}p_{\theta}\left(\bm{z}_{t-1}|\bm{z}_{t}\right),\end{split}$ (5)
@@
[…]
Figure 3: Samples from unconditional generation. Our method produces high-quality and diverse shapes. We also calculate their average CD to each object in the training set to confirm that our model is capable of producing unique shapes.
@@ 3.3 Conditional Shape Generation
After training, we offer support for four types of conditional shape generation: category-conditioned shape generation, sketch-conditioned shape generation, single-view conditioned shape reconstruction, and text-guided shape generation.
@@ 3.4 Training Details
[…]
Figure 4: Qualitative results of Category Conditional Generation. Our method generates different categories of detailed 3D shapes with high-quality and diversity.
[…]
@@ 4 Experiments
We conduct extensive qualitative and quantitative experiments to demonstrate the generation capability of our method on five tasks: unconditional generation (Sec.4.2), category-conditioned shape generation (Sec.4.3.1), sketch-conditioned shape generation (Sec.4.3.2), single-view 3D reconstruction (Sec.4.3.3), and text-guided generation (Sec.4.3.4). We conduct ablation studies to evaluate our method in Sec.4.4.
@@ 4.2 Unconditional Shape Generation
For unconditional generation, we compare our method with SDFusion [12] and LAS-Diffusion [76], which recently demonstrated state-of-the-art results on 3D shape generation tasks. We follow Yang et al. [73] to use minimum matching distance (MMD), coverage (COV), and 1-nearest neighbor accuracy (1-NNA) based on the Chamfer distance  [49] for evaluation metrics. MMD measures quality, COV measures diversity, and 1-NNA measures the similarity of the ground truth and generated distributions. Lower MMD, higher COV, 1-NNA that has a smaller difference to $50\%$ mean better quality. We generate the same number of samples as the test set for evaluation.
As shown in Tab. 1, our method outperforms baselines in these metrics. Our method gets significantly better scores in the MMD metric which means our method can generate more realistic and detailed shapes. As the qualitative results are shown in Fig. 3, our method can generate high-quality and diverse clothes. We visualize the generated samples according to the average distance between the generations and each object in the training set.
@@ 4.3 Conditional Shape Generation
4.3.1 Category Conditional Generation
In this section, we demonstrate category conditional generation experiments in 9 categories of no sleeve shirt, short sleeve shirt, long sleeve shirt, no sleeve dress, short sleeve dress, long sleeve dress, dress, short pants, and long pants objects in Deepfashion3D dataset by conditioning in category information. We use the same evaluation metrics as Tab. 1 to evaluate the generated results. The quantitative comparison of shirt, dress, pants objects are in Tab. 2. As shown in Tab. 2, our method outperforms SDFusion in all metrics. We also visualize generated 3D shapes in Fig. 4. Our method generates different categories of detailed 3D shapes with high-quality and diversity.
@@ 4.3.1 Category Conditional Generation
[…]
4.3.2 Sketch-guided Shape Generation
For sketch-conditioned shape generation, we follow Zheng et al. [76] to generate 2D sketches for each shape in the dataset by rendering its shading images from different views and extracting their edges via the Canny edge detector [3]. The Chamfer distance (CD), Earth mover’s distance (EMD), and the Voxel-IOU are used as evaluation metrics. We compare our method with two state-of-the-art methods for 3D reconstruction and generation, SDFusion [12] and LAS-Diffusion [76]. Quantitatively, as shown in Tab. 3, our method outperforms other methods on all metrics. As shown in Fig. 5, SDFusion and LAS-Diffusion generate unrealistic results in open surfaces as the drawback of SDF representations, while our method generates high-quality and detailed results aligned with the input sketch.
@@ 4.3.2 Sketch-guided Shape Generation
4.3.3 Single-view Shape Reconstruction
@@
4.3.4 Text-guided Shape Generation
@@
4.4 AutoEncoder Structure Ablation
In this section, we investigate the effectiveness of our encoder design for shape embedding. Specifically, we compare discrete volume-based representation the same as SDFusion [12] for UDFs with ours where UDFs have continuous and accessible gradient fields. We test by overfitting the models with corresponding settings to a garment. As shown in Fig. 8, our model faithfully captures the input shape, yielding a lower Chamfer Distance (CD) value of $0.46$, whereas the grid-based model struggles to accurately reconstruct the surface with a higher CD value $8.37$ (the unit of CD is $10^{-3}$). Note that the zero-level set is not extracted due to unreliable gradient estimation. We extract an $\epsilon$-iso surface with $\epsilon=0.02$ for comparison.
@@ 4.4 AutoEncoder Structure Ablation
Figure 8: Influence of different UDF encoding approaches.

### 23 Mar 2024 v2

# Surf-D: Generating High-Quality Surfaces of Arbitrary Topologies Using Diffusion Models

@@ Abstract
We present Surf-D, a novel method for generating high-quality 3D shapes as Surfaces with arbitrary topologies using Diffusion models. Previous methods explored shape generation with different representations and they suffer from limited topologies and poor geometry details. To generate high-quality surfaces of arbitrary topologies, we use the Unsigned Distance Field (UDF) as our surface representation to accommodate arbitrary topologies. Furthermore, we propose a new pipeline that employs a point-based AutoEncoder to learn a compact and continuous latent space for accurately encoding UDF and support high-resolution mesh extraction. We further show that our new pipeline significantly outperforms the prior approaches to learning the distance fields, such as the grid-based AutoEncoder, which is not scalable and incapable of learning accurate UDF. In addition, we adopt a curriculum learning strategy to efficiently embed various surfaces. With the pretrained shape latent space, we employ a latent diffusion model to acquire the distribution of various shapes. Extensive experiments are presented on using Surf-D for unconditional generation, category conditional generation, image conditional generation, and text-to-shape tasks. The experiments demonstrate the superior performance of Surf-D in shape generation across multiple modalities as conditions. Visit our project page at https://yzmblog.github.io/projects/SurfD/.
Keywords:
Diffusion Model Shape Generation Unsigned Distance Fields Generative Model
1 Introduction
[…]
@@ 1 Introduction
Generating a diverse array of 3D shapes using various conditions is an important problem in computer vision and graphics, given its wide applications, such as gaming, filmmaking, and the metaverse. Recent advancements for 3D shape generation [5, 1, 41, 30, 20, 33, 39] aim to make this process more accessible.
@@
The diverse interplay of topology and intricate geometry gives rise to countless 3D shapes in our world. However, existing generation works still suffer from limited topologies and the lack of geometric details. Previous methods present shapes using point clouds [1, 41], voxel representations [58, 72], and signed distance fields (SDFs) [11, 45].
[…]
Figure 2: Comparison of learning discrete grid-based field [12] and continuous field (ours) for UDFs (top) and corresponding gradient errors (bottom).The darker the color, the smaller the values.
@@
However, point cloud and voxel representations encounter difficulties in faithfully extracting surfaces. Implicit representations, particularly SDFs, excel in describing closed surfaces but face challenges with open surfaces and complex topologies. Despite attempts to enhance them, e.g. 3PSDF [10] and TSDF [61], the representations remain reliant on the underlying definition of SDF. G-Shell [35] is proposed to represent open surfaces but still falters with non-manifold structures and computational constraints.
Previous 3D shape generation methods based on learning a distance field include SDFusion [12] and its variants [28, 76]. These methods consider only SDF and all use a discrete 3D feature grid to approximate the SDF field.
This grid-based approach is fundamentally limited and unsuitable for learning high-quality UDF because of the following drawbacks: (i) it does not provide access to accurate and stable gradient of the distance field, which is essential for accurately reconstructing the surface encoded by the UDF; (ii) limited resolution of the generated grid; (iii) expensive memory cost for the grid representation.
With these disadvantages, the grid-based framework cannot be used for learning UDF to generate high-quality surfaces of arbitrary topologies, as confirmed by our experiments. Briefly, Fig. 2 shows that the UDF generated using the grid-based method [12] suffers from unstable, as shown in the gradient field error map [^1] We calculate the numerical gradients of UDF following Neuralangelo [30] for visualizing the error map and meshing. in Fig. 2 (a) bottom, in comparison with our results in Fig. 2 (b) bottom. These unreliable gradients will in turn lead to low-quality surface meshes.
To tackle the aforementioned problems, We propose Surf-D, the first generative latent diffusion model [53] employing Unsigned Distance Fields (UDF) which is a general shape representation for any 3D shape, for versatile shape generation with high-quality surfaces of arbitrary topologies. Unlike previous discrete grid-based shape embedding methods [12, 28, 76], Surf-D utilizes a point-based AutoEncoder to effectively capture UDFs by compressing the underlying 3D surfaces into compact low-dimensional latent codes and learns Continuous Distance Fields, which allows for a more accurate and robust gradient field computation, as evidenced by the gradient error map in Fig. 2 (b) bottom. Compared with existing discrete-grid-based approaches, our Surf-D enjoys several advantages: (i) reliable gradient query for each input point which supports accurate mesh extraction [21]; (ii) capturing UDFs without a restriction by resolution, as it learns a continuous field; (iii) high computational efficiency; See detailed experiments and analysis between our approach and discrete grid-based shape embedding in Sec. 4.5.
@@
Moreover, to learn the compact shape latent space efficiently and effectively on diverse surfaces, we employ a curriculum learning strategy [2] to train our model in an easy-to-hard sampling order, i.e. progressively extending the current training dataset during training. This strategy promotes our model to develop a diverse latent space, avoiding entrapment into generating mean shape, i.e. model collapse. As a result, our model is not only much more lightweight and learning-efficient but also feasible to convert high-quality generated implicit shapes to explicit meshes.
@@
    We conduct latent diffusion and devise a point-based AutoEncoder that enables effective gradient querying for given input points in the field, achieving high-resolution and detail-preserving surface generation. To enhance surface learning, we implement a curriculum learning scheme.
@@ 2 Related Work
Tremendous efforts have been made for 3D shape generation [5, 1, 41, 4, 20, 34, 7], given its broad applications, employing various shape representations such as voxel grids [70, 68], point clouds [41, 64, 78], and implicit neural fields [48, 17, 12, 24, 40]. Additionally, various methods have been developed to guide the synthesis with text prompt [55, 18, 37, 9, 36], single-view image [33, 39, 32, 31], sketch [71, 76] and sparse control anchors [42, 13]. In the early stages, voxel-based generation methods [70, 68, 51] can synthesize various coarse shapes. However, the voxel representation can not scale to high-resolution generation [77, 76], due to its cubic space complexity. Utilizing point clouds [41, 64, 78], which are compact and readily available, for 3D object representation can achieve arbitrary shape generation with their well-suitability to learn geometry and texture information [46, 71]. However, it requires complex post-processing to faithfully convert the synthesized point clouds into meshes for real-world applications [49].
@@ 3.1 Point-based AutoEncoder
Curriculum Learning Strategy
@@
Curriculum training involves measuring performance and establishing an easy-to-hard scheduler. In this paper, we use a simple yet effective performance metric, which is the training loss $\mathcal{L}_{latent}$. We select samples with lower losses as incoming training data in each iteration. This continues until all data is included. Specifically, we first randomly sample $k$ shapes from the dataset, then gradually add new difficult samples according to their loss values, resulting in an easy-to-hard gradual learning process. The pseudo-code can be expressed as Algo. 1. The influence of the proposed learning strategy is included in the Appendix.
Algorithm 1 Curriculum Learning Strategy for Shape Embedding
1:$M$ – our model;
2:$\mathcal{D}$ – the whole dataset;
3:$\mathcal{D}^{*}$ – current training dataset;
4:$n$,$s$,$k$ – the number of training epochs, switching epochs, and selected samples;
$\mathcal{L}$ – Loss calculation;
6:$\mathop{\mathrm{arg\,kmin}}$ – Find the subset containing $k$ elements with minimal scores;
7:
8:$\mathcal{D}^{*}\leftarrow$random select $k$ samples from $\mathcal{D}$
9:for $i=1$ to $n$ do
10:     if ($t\%s$ == 00) and (not $\mathcal{D}^{*}$ == $\mathcal{D}$then
11:         $\mathcal{D}^{*}\leftarrow\mathcal{D}^{*}\cup\mathop{\mathrm{arg\,kmin}}\limits_{x\in\mathcal{D}-\mathcal{D}^{*}}{\mathcal{L}(M,x)}$
12:     end if
13:     $M\leftarrow train(M,E^{*},C)$
14:end for
@@ 3.2 Latent Diffusion Model for UDF
$\begin{split}q\left(\bm{z}_{0:T}\right)&=q\left(\bm{z}_{0}\right)\prod_{t=1}^{T}q\left(\bm{z}_{t}|\bm{z}_{t-1}\right),\\ p_{\theta}\left(\bm{z}_{0:T}\right)&=p\left(\bm{z}_{T}\right)\prod_{t=1}^{T}p_{\theta}\left(\bm{z}_{t-1}|\bm{z}_{t}\right),\end{split}$ (5)
@@ 3.3 Conditional Shape Generation
We evaluate Surf-D on four conditional shape generation tasks: category-conditioned shape generation, sketch-conditioned shape generation, single-view conditioned shape reconstruction, and text-guided shape generation.
@@ 4 Experiments
We conduct extensive qualitative and quantitative experiments to demonstrate the generation capability of Surf-D on five tasks: unconditional generation in Sec. 4.2, category-conditioned shape generation, sketch-conditioned shape generation, single-view 3D reconstruction, and text-guided generation in Sec. 4.3. We evaluate different AutoEncoder designs in Sec. 4.5 and latent space interpolation in Sec. 4.6.
@@ 4.1 Dataset
[…]
Figure 4: Samples from unconditional generation. Our method produces high-quality and diverse shapes. We also calculate their average CD for each object in the training set to confirm that our model is capable of producing unique shapes.
@@ 4.2 Unconditional Shape Generation
For unconditional generation, we compare our method with SDFusion [12] and LAS-Diffusion [76], which recently demonstrated state-of-the-art results on 3D shape generation tasks. We follow Yang et al. [74] to use minimum matching distance (MMD), coverage (COV), and 1-nearest neighbor accuracy (1-NNA) based on the Chamfer distance  [50] for evaluation. MMD measures quality, COV measures diversity, and 1-NNA measures the similarity of the ground truth and generated distributions. Lower MMD, higher COV, 1-NNA that has a smaller difference to $50\%$ mean better quality. We generate the same number of samples as the test set for evaluation.
As shown in Tab. 1, our method outperforms baselines in these metrics. Our method gets significantly better scores in the MMD metric which means our method can generate more realistic and detailed shapes. As shown in Fig. 4, our method can generate high-quality and diverse clothes. We visualize the generated samples according to the average distance between the generations and each object in the training set.
@@
Figure 5: Qualitative results of Category Conditional Generation. Our method generates different categories of detailed 3D shapes with high quality and diversity.
@@ 4.3 Conditional Shape Generation
Category Conditional Generation
In this section, we demonstrate category conditional generation experiments in 9 categories of no sleeve shirt, short sleeve shirt, long sleeve shirt, no sleeve dress, short sleeve dress, long sleeve dress, dress, short pants, and long pants objects in Deepfashion3D dataset by conditioning in category information. We use the same evaluation metrics as Tab. 1 to evaluate the generated results. The quantitative comparison of shirt, dress, pants objects are in Tab. 2. As shown in Tab. 2, our method outperforms SDFusion in all metrics. We also visualize generated 3D shapes in Fig. 5. Our method generates different categories of detailed 3D shapes with high quality and diversity.
Sketch-guided Shape Generation
@@
For sketch-conditioned shape generation, we follow Zheng et al. [76] to generate 2D sketches for each shape in the dataset by rendering its shading images from different views and extracting their edges via the Canny edge detector [3]. The Chamfer distance (CD), Earth mover’s distance (EMD), and the Voxel-IOU are used as evaluation metrics. We compare our
[…]
@@
method with two state-of-the-art methods for 3D reconstruction and generation, SDFusion [12] and LAS-Diffusion [76]. Quantitatively, as shown in Tab. 3, our method outperforms other methods on all metrics. As shown in Fig. 6, SDFusion and LAS-Diffusion generate unrealistic results in open surfaces as the drawback of SDF representations, while our method generates high-quality and detailed results aligned with the input sketch.
@@
Single-view Shape Reconstruction
@@
[…]
Text-guided Shape Generation
[…]
@@
Figure 9: Garment generation for virtual try-on . The generated garments can seamlessly integrate into a simulation pipeline for virtual try-on.
4.4 Shape for Simulation
Furthermore, we show that the generated clothing can be easily integrated into a simulation platform [^1] https://www.marvelousdesigner.com/ for virtual try-on, which is shown in Fig. 9. We use the motion sampled from AMASS [43] mocap dataset. Please refer to the video on our webpage in the supplementary material for the visualization. These compelling results validate our high-quality 3D asset generation.
4.5 Grid-based Distance Fields V.S. Continuous Distance Fields
@@ 4.4 Shape for Simulation
Figure 10: Comparison of different UDF encoding approaches on shape generation. Compared with the grid-based manner, our method produces high-quality results by learning continuous distance fields.
Table 4: Memory cost comparison.
  Methods Resolution Model Size
  SDFusion [12] $64\times 64\times 64$ $4.29$G
Diffusion-SDF [28] $64\times 64\times 64$ $5.86$G
LAS-Diffusion [76] $128\times 128\times 128$ $859$M
  Ours $\mathbf{512\times 512\times 512}$ $\mathbf{564}$M
In this section, we investigate the effectiveness of our point-based AutoEncoder design for shape embedding, in comparison with the grid-based approach. Specifically, we apply the discrete grid-based representation (as used in SDFusion [12]) to learning UDF, and then compare it with with Surf-D. We test by overfitting the models with corresponding settings to garments and close watertight objects(chairs and tables). As shown in Fig. 10, our model faithfully reconstructs the input shapes with high quality, yielding a lower Chamfer Distance (CD) value, whereas the grid-based model struggles to accurately reconstruct the surface with a higher CD value (the unit of CD is $10^{-3}$). The grid-based representation can not capture implicit fields in detail compared to learning continuous distance fields. A similar observation is also reflected in SDF [48], which is consistent with our observations. Besides, we also compare the resolution and memory cost between our Surf-D with other SOTA methods. As shown in Tab. 4, our method achieves higher resolution with lower model size as we learn more compact latent codes compared with other methods. More details of comparison are provided in the Appendix.